A typist makes an average of 4 typo errors per page that she types. If the number of errors per each typed page can be modeled according to the Poisson distribution, what is the probability that she will make more than 3 typo errors in the next page? Express your answer in decimals and round to two significant places after the decimal. For example 0.1678 should be entered as 0.17.
Answer #1

X ~ Poisson ()

Where = 4 errors per page.

Poisson probability distribution is

P(X) = e- * X / X!

P(X > 3) = 1 - P( X <= 3)

= 1 - [ P( X = 0) + P (X = 1) + P( X = 2) + P( X = 3) ]

= 1 - [ e-4 + e-4 * 4 + e-4​​​​​​​ * 42 / 2 ! + e-4​​​​​​​ * 43 / 3! ]

= 0.57